# Topic-Based Watermarks for Large Language Models

### 02 Mar 2026 v5

@@ Abstract
The indistinguishability of large language model (LLM) output from human-authored content poses significant challenges, raising concerns about potential misuse of AI-generated text and its influence on future model training. Watermarking algorithms offer a viable solution by embedding detectable signatures into generated text. However, existing watermarking methods often involve trade-offs among attack robustness, generation quality, and additional overhead such as specialized frameworks or complex integrations. We propose a lightweight, topic-guided watermarking scheme for LLMs that partitions the vocabulary into topic-aligned token subsets. Given an input prompt, the scheme selects a relevant topic-specific token list, effectively “green-listing” semantically aligned tokens to embed robust marks while preserving fluency and coherence. Experimental results across multiple LLMs and state-of-the-art benchmarks demonstrate that our method achieves text quality comparable to industry-leading systems and simultaneously improves watermark robustness against paraphrasing and lexical perturbation attacks, with minimal performance overhead. Our approach avoids reliance on additional mechanisms beyond standard text generation pipelines, enabling straightforward adoption and suggesting a practical path toward globally consistent watermarking of AI-generated content.
@@
Alexander Nemecek1, Yuzhou Jiang2, Erman Ayday1, 1Case Western Reserve University, 2Meta Platforms, Inc., Correspondence: ajn98@case.edu
@@ 1 Introduction
Researchers have explored watermarking as a complementary or alternative solution. Rather than detecting AI-authored text post hoc, watermarking algorithms embed a detectable signature into text during the generation process Kirchenbauer et al. (2023); Scott Aaronson (2023). By introducing controlled token-level modifications, watermarks can remain identifiable even after moderate transformations. However, existing approaches face trade-offs where methods prioritizing computational efficiency lack robustness to paraphrasing Dathathri et al. (2024), while more robust techniques require costly architectural modifications or multiple inference passes that degrade practical deployability Huo et al. (2024); Kuditipudi et al. (2024).
@@ 2 Related Work
Traditional AI-generated text detection relied on post hoc classifiers trained to differentiate human and machine writing, but these often fail under adversarial transformations like paraphrasing. To address this, watermarking techniques have gained prominence as a more robust alternative, categorized into post-processing and generation methods. Post-processing watermarking modifies generated text to embed hidden patterns but remains vulnerable to edits that erase markers Sato et al. (2023). Generation watermarking embeds signals during token generation, ensuring greater resilience against text modifications.
@@ 3.2 Threat Model
We assume a standard black-box adversary with access only to LLM outputs, seeking to remove watermarks while preserving semantic fidelity. Following Kirchenbauer et al. (2023); Zhang et al. (2024), the adversary cannot access model parameters, watermarking keys, or topic-partition logic. While the detection algorithm is public, specific parameters (thresholds, embeddings) remain private. We evaluate robustness against two primary attack surfaces: (1) text degradation via token-level perturbations (e.g., substitutions) that disrupt the watermark signal at the cost of fluency; and (2) semantic rephrasing using external models for paraphrasing or summarization to rewrite content while maintaining meaning. Detailed adversary assumptions and excluded attack vectors (e.g., statistical comparison) are provided in B.
@@ 4.1 Token-to-Topic Mappings
The hyperparameter $\tau$ controls the granularity of semantic alignment and comprehensive topic coverage where a higher $\tau$ enforces stronger coherence but increases the proportion of tokens allocated via the round-robin mechanism. Although we consider only four broad topics in our main evaluations, the same procedure naturally extends to a larger number of topics for more fine-grained coverage as the vocabulary $V$ increases. We use a general-purpose sentence model (all-MiniLM-L6-v2) for this implementation as a lightweight approach; however, any semantic embedding framework can be substituted, allowing practitioners to tailor the mapping for domain-specific or resource-constrained environments. Figure 1 illustrates this token-to-topic assignment process, with the detailed implementation procedure for provided in Algorithm 1 (A).
@@ 5.1 Experimental Setup
Watermarking Comparisons. We compare TBW against several baselines (see §2 for details): No Watermark (standard decoding), KGW Kirchenbauer et al. (2023), DipMark (DiP) Wu et al. (2024), Unigram Zhao et al. (2024), SynthID-Text (SynthID) Dathathri et al. (2024), SIR Liu et al. (2024), and EXP, EXP-Edit, ITS-Edit Kuditipudi et al. (2024). We use the MarkLLM library Pan et al. (2024) implementations with parameter settings from the original papers to ensure fair comparison (details in D).
Implementation Details: Our TBW uses KeyBert Grootendorst (2020) for topic extraction alongside a partition of topic-aligned tokens (§4.1). We employ a predefined set of four generalized topics, {animals, technology, sports, medicine}, which we found to be sufficiently generic to cover large vocabulary partitions while avoiding over-specialization. In our setup, four topics correspond to an effective green-list fraction of $0.25$, which is close to the $\gamma=0.2$ used in KGW’s experiments, while still maintaining broad topic coverage. Topic assignment is performed using a sentence embedding model (all-MiniLM-L6-v2) with a similarity threshold of $\tau=0.7$, though any semantic embedding framework can be substituted. We also study scalability beyond four topics by varying $K\in\{4,8,16,32\}$; as $K$ increases the watermark signal attenuates as expected yet remains strong, and text quality stays flat (see F.4 for details).
@@ 5.3 Robustness
ROC–AUC Best F1 Score
@@
Table 1: ROC–AUC and Best F1 scores for watermarking approaches under no attack, PEGASUS paraphrasing, and DIPPER paraphrasing. Best results are in bold; second-best results are underlined.
@@
5.3.1 Text Degradation
@@
5.3.2 Semantic Paraphrasing
We first report ROC-AUC and Best F1 scores in Table 1. Without attack, all methods reliably separate watermarked from non-watermarked text. Under paraphrasing, however, DIPPER and PEGASUS substantially degrade the performance of SynthID and DiP. TBW and Unigram retain the highest ROC-AUC and Best F1 scores across both attack scenarios, with TBW often matching Unigram and outperforming all other baselines. Full ROC-AUC curves are included in E.6.
In practical applications, maintaining a high True Positive Rate (TPR) at consistently low False Positive Rates (FPR) ensure non-watermarked texts are not misclassified as watermarked. We therefore report TPR@1%FPR and TPR@10%FPR for OPT-6.7B and Gemma-7B in Table 2, with additional human-text FPR analysis in E.7.
5.4 Detection Analysis
The maximum $z$-score detection method delivers near-perfect results, achieving 99.6% and 100% detection rates for OPT-6.7B and GEMMA-7B, respectively, with the highest mean $z$-scores and minimal variance. This performance stems from its independence from topic extraction by directly evaluating all topic partitions and selecting the highest $z$-score, it avoids the alignment errors that constrain topic-dependent approaches. We confirm this with 500 watermarked and 500 non-watermarked samples with detailed performance metrics provided in E.9. These results connect our oracle-condition evaluations in §5.3 with practical, topic-agnostic detection, showing that maximum $z$-score preserves robustness while providing tunable separation adaptable to different operational constraints.
@@ 5.5 Efficiency
We measure the computational overhead imposed by each watermarking method on OPT-6.7B and OPT-2.7B, generating sequences of lengths {100, 200, 300, 400, 500}. For each token-length, we record the average generation time over 10 samples from C4. TBW introduces negligible overhead compared to non-watermarked generation, matching lightweight methods such as KGW and SynthID while EXP-based approaches incur substantial costs from multiple decoding passes. Full timing results are provided in E.4.
@@ 6 Conclusion
Watermarking has emerged as a critical tool for mitigating risks in large language models, yet existing schemes face a persistent tradeoff where lightweight methods such as SynthID-Text or KGW offer efficiency and text quality but collapse under paraphrasing, while heavyweight multi-pass methods deliver robustness at the cost of fluency and practicality. We presented Topic-Based Watermarking (TBW), which addresses this divide by embedding semantically aligned signals through topic-driven token partitioning. TBW preserves the speed and fluency of lightweight approaches while achieving robustness closer to heavyweight schemes. Our evaluation confirms that TBW provides strong detection accuracy, resilience to lexical and semantic attacks, and negligible runtime overhead. By overcoming this tradeoff, TBW offers a practical path toward deployable watermarking.
@@
Beyond TBW’s specific implementation, a limitation it shares with other list-based watermarking schemes, including KGW Kirchenbauer et al. (2023) and SynthID-Text Dathathri et al. (2024), is reduced detection reliability on shorter text sequences. Our primary evaluations use 200-token generations, which we consider relatively short for watermark detection. The statistical nature of $z$-score-based detection requires sufficient tokens to distinguish the watermark signal from random variation; as text length decreases, the gap between watermarked and non-watermarked $z$-score distributions narrows, increasing the likelihood of misclassification. We hypothesize that detection performance would degrade on shorter sequences, a constraint inherent to the underlying statistical framework.
@@
Watermarking techniques inherently carry dual-use risks. While watermarking enables provenance tracking to combat misuse, it could potentially be exploited for punitive enforcement based on probabilistic signals. Adversarial actors may further weaponize watermarking to falsely attribute content to specific organizations or models. We additionally acknowledge residual concerns regarding false positives, which could lead to wrongful accusations. Accordingly, we recommend deployment with conservative thresholds, transparent documentation of error rates, and appeals processes for affected users, calibrated to the specific deployment context.
@@
      False Positive Rate (FPR) on Human-Written Text ........................................................................................................................................................................E.7
@@
      Detection Attacks ........................................................................................................................................................................E.8
@@
      Maximum $z$-Score Detection: Detailed Analysis ........................................................................................................................................................................E.9
@@ Appendix B Extended Threat Model
Excluded and Impractical Attacks. We further assume the adversary cannot mount impractical strategies that rely on excessive querying or access to paired datasets (watermarked vs. non-watermarked versions of the same prompt). Statistical comparison or supervised classification attacks are excluded from our primary model as they require assumptions that are rarely feasible for typical end-users. We provide an empirical evaluation of these “detection-based” attacks in E.8 for completeness, though they fall outside our primary security focus.

### 15 Apr 2026 v6

@@ Abstract
The indistinguishability of large language model (LLM) output from human-authored content poses significant challenges, raising concerns about potential misuse of AI-generated text and its influence on future model training. Watermarking algorithms offer a viable solution by embedding detectable signatures into generated text. However, existing watermarking methods often involve trade-offs among attack robustness, generation quality, and additional overhead such as specialized frameworks or complex integrations. We propose a lightweight, topic-guided watermarking scheme for LLMs that partitions the vocabulary into topic-aligned token subsets. Given an input prompt, the scheme selects a relevant topic-specific token list, effectively “green-listing” semantically aligned tokens to embed robust marks while preserving fluency and coherence. Experimental results across multiple LLMs and state-of-the-art benchmarks demonstrate that our method achieves text quality comparable to industry-leading systems and simultaneously improves watermark robustness against paraphrasing and lexical perturbation attacks, with minimal performance overhead. Our approach avoids reliance on additional mechanisms beyond standard text generation pipelines, enabling straightforward adoption and suggesting a practical path toward globally consistent watermarking of AI-generated content. Our code is available at https://github.com/ANCP2021/Topic-Based-Watermarks.
@@
Alexander Nemecek1thanks: Correspondence: ajn98@case.edu., Yuzhou Jiang2, Erman Ayday1 1Case Western Reserve University, 2Meta Platforms, Inc.
@@ 1 Introduction
Researchers have explored watermarking as a complementary or alternative solution. Rather than detecting AI-authored text post-hoc, watermarking algorithms embed a detectable signature into text during the generation process Kirchenbauer et al. (2023); Scott Aaronson (2023). By introducing controlled token-level modifications, watermarks can remain identifiable even after moderate transformations. However, existing approaches face trade-offs where methods prioritizing computational efficiency lack robustness to paraphrasing Dathathri et al. (2024), while more robust techniques require costly architectural modifications or multiple inference passes that degrade practical deployability Huo et al. (2024); Kuditipudi et al. (2024).
@@ 2 Related Work
Traditional AI-generated text detection relied on post-hoc classifiers trained to differentiate human and machine writing, but these often fail under adversarial transformations like paraphrasing. To address this, watermarking techniques have gained prominence as a more robust alternative, categorized into post-processing and generation methods. Post-processing watermarking modifies generated text to embed hidden patterns but remains vulnerable to edits that erase markers Sato et al. (2023). Generation watermarking embeds signals during token generation, ensuring greater resilience against text modifications.
@@ 3.2 Threat Model
We assume a standard black-box adversary with access only to LLM outputs, seeking to remove watermarks while preserving semantic fidelity. Following Kirchenbauer et al. (2023); Zhang et al. (2024), the adversary cannot access model parameters, watermarking keys, or topic-partition logic. While the detection algorithm is public, specific parameters (thresholds, embeddings) remain private. We evaluate robustness against two primary attack surfaces: (1) text degradation via token-level perturbations (e.g., substitutions) that disrupt the watermark signal at the cost of fluency; and (2) semantic rephrasing using external models for paraphrasing or summarization to rewrite content while maintaining meaning.
We additionally consider (3) partition recovery attacks, where an adversary attempts to reconstruct the topic-aligned green lists by exploiting the semantic structure of TBW’s vocabulary partitioning. Prior work has demonstrated such attacks against KGW-style schemes through repeated positional querying Jovanović et al. (2024). While TBW’s topic-aligned partitions introduce semantic co-grouping that could appear more predictable than random assignment, the full partition remains unrecoverable without the private random seed used during round-robin distribution of residual tokens (§4). Detailed adversary assumptions, excluded attack vectors (e.g., statistical comparison), detection-based attacks, and the full partition recovery analysis are provided in B.
@@ 4.1 Token-to-Topic Mappings
The hyperparameter $\tau$ controls the granularity of semantic alignment and comprehensive topic coverage where a higher $\tau$ enforces stronger coherence but increases the proportion of tokens allocated via the round-robin mechanism. Although we consider only four broad topics in our main evaluations, the same procedure naturally extends to a larger number of topics for more fine-grained coverage as the vocabulary $V$ increases. We use a general-purpose sentence model (all-MiniLM-L6-v2) for this implementation as a lightweight approach; however, any semantic embedding framework can be substituted, allowing practitioners to tailor the mapping for domain-specific or resource-constrained environments. Figure 1 illustrates this token-to-topic assignment process, with the detailed implementation procedure provided in Algorithm 1 (A).
@@ 5.1 Experimental Setup
Watermarking Comparisons. We compare TBW against several baselines (see §2 for details): No Watermark (standard decoding), KGW Kirchenbauer et al. (2023), DiPMark (DiP) Wu et al. (2024), Unigram Zhao et al. (2024), SynthID-Text (SynthID) Dathathri et al. (2024), SIR Liu et al. (2024), and EXP, EXP-Edit, ITS-Edit Kuditipudi et al. (2024). We use the MarkLLM library Pan et al. (2024) implementations with parameter settings from the original papers to ensure fair comparison (details in D). We additionally compare TBW against semantic watermarking schemes that operate via post-processing rather than in-generation biasing, specifically PostMark Chang et al. (2024) and k-SemStamp Hou et al. (2024b); these results are reported separately in E.6 due to their different generation pipelines.
Implementation Details: Our TBW uses KeyBERT Grootendorst (2020) for topic extraction alongside a partition of topic-aligned tokens (§4.1). We employ a predefined set of four generalized topics, {animals, technology, sports, medicine}, which we found to be sufficiently generic to cover large vocabulary partitions while avoiding over-specialization. In our setup, four topics correspond to an effective green-list fraction of $0.25$, which is close to the $\gamma=0.2$ used in KGW’s experiments, while still maintaining broad topic coverage. Topic assignment is performed using a sentence embedding model (all-MiniLM-L6-v2) with a similarity threshold of $\tau=0.7$, though any semantic embedding framework can be substituted. We also study scalability beyond four topics by varying $K\in\{4,8,16,32\}$; as $K$ increases the watermark signal attenuates as expected yet remains strong, and text quality stays flat (see 5.6 and F.4).
@@ 5.3 Robustness
5.3.1 Text Degradation
5.3.2 Semantic Paraphrasing
We first report ROC-AUC and Best F1 scores in Table 1. Without attack, all methods reliably separate watermarked from non-watermarked text. Under paraphrasing, however, DIPPER and PEGASUS substantially degrade the performance of SynthID and DiP. TBW and Unigram retain the highest ROC-AUC and Best F1 scores across both attack scenarios, with TBW often matching Unigram and outperforming all other baselines. Full ROC-AUC curves are included in E.7.
In practical applications, maintaining a high True Positive Rate (TPR) at consistently low False Positive Rates (FPR) ensure non-watermarked texts are not misclassified as watermarked. We therefore report TPR@1%FPR and TPR@10%FPR for OPT-6.7B and Gemma-7B in Table 2, with additional human-text FPR analysis in E.8.
ROC-AUC Best F1 Score
@@ 5.3.2 Semantic Paraphrasing
Table 1: ROC-AUC and Best F1 scores for watermarking approaches under no attack, PEGASUS paraphrasing, and DIPPER paraphrasing. Best results are in bold; second-best results are underlined.
@@
5.4 Detection Analysis
The maximum $z$-score detection method delivers near-perfect results, achieving 99.6% and 100% detection rates for OPT-6.7B and GEMMA-7B, respectively, with the highest mean $z$-scores and minimal variance. This performance stems from its independence from topic extraction by directly evaluating all topic partitions and selecting the highest $z$-score, it avoids the alignment errors that constrain topic-dependent approaches. We confirm this with 500 watermarked and 500 non-watermarked samples with detailed performance metrics provided in E.10. These results connect our oracle-condition evaluations in §5.3 with practical, topic-agnostic detection, showing that maximum $z$-score preserves robustness while providing tunable separation adaptable to different operational constraints.
@@ 5.5 Efficiency
We measure the computational overhead imposed by each watermarking method on OPT-6.7B, generating sequences of lengths {100, 200, 300, 400, 500}. For each token-length, we record the average generation time over 10 samples from C4. Figure 4 shows that TBW introduces negligible overhead compared to non-watermarked generation, matching lightweight methods such as KGW and SynthID. In contrast, EXP-Edit requires multiple re-ranking passes and SIR incurs additional complexity, resulting in noticeably higher generation times. We observe consistent trends on OPT-2.7B (see E.4), confirming that TBW’s efficiency advantage holds across model scales.
5.6 Topic Scalability
While our main experiments use $K=4$ topics, practitioners may require finer-grained coverage. We evaluate TBW with $K\in\{4,8,16,32\}$ on Gemma-7B, whose larger vocabulary is more representative of deployed LLMs. We apply $\delta=2.0$ and $\tau=0.5$, relaxed relative to the main experiments to compensate for smaller per-topic coverage at larger $K$. Detection uses the maximum $z$-score scheme. Full experimental details, topic inventories, and additional visualizations are provided in F.4.
Figure 5 reports detection strength as a function of $K$. The $z$-score decreases gracefully from ${\sim}11$ at $K=4$ to ${\sim}7$ at $K=32$, yet remains comparable to KGW and Unigram baselines at equivalent watermark strength. Text quality (BERTScore F1) remains flat across all $K$ values (see F.4), confirming that increased topic granularity does not degrade fluency or semantics. Table 4 reports detection runtime, which scales approximately linearly with $K$ but remains tractable; because detection is performed offline by the model owner, this overhead does not impact user-facing latency.
[…]
Number of Topics ($K$) Detection Time (s)
4 $2.002\pm 0.041$
8 $2.910\pm 0.047$
16 $3.833\pm 0.046$
32 $6.565\pm 0.066$
@@ 6 Conclusion
Watermarking has emerged as a critical tool for mitigating risks in large language models, yet existing schemes face a persistent trade-off where lightweight methods such as SynthID-Text or KGW offer efficiency and text quality but collapse under paraphrasing, while heavyweight multi-pass methods deliver robustness at the cost of fluency and practicality. We presented Topic-Based Watermarking (TBW), which addresses this divide by embedding semantically aligned signals through topic-driven token partitioning. TBW preserves the speed and fluency of lightweight approaches while achieving robustness closer to heavyweight schemes. Our evaluation confirms that TBW provides strong detection accuracy, resilience to lexical and semantic attacks, and negligible runtime overhead. By overcoming this trade-off, TBW offers a practical path toward deployable watermarking.
@@
Beyond TBW’s specific implementation, a limitation it shares with other list-based watermarking schemes, including KGW Kirchenbauer et al. (2023) and SynthID-Text Dathathri et al. (2024), is reduced detection reliability on shorter text sequences. Our primary evaluations use 200-token generations, which we consider relatively short for watermark detection. The statistical nature of $z$-score-based detection requires sufficient tokens to distinguish the watermark signal from random variation; as text length decreases, the gap between watermarked and non-watermarked $z$-score distributions narrows, increasing the likelihood of misclassification. We hypothesize that detection performance would degrade on shorter sequences, a constraint inherent to the underlying framework.
@@
Watermarking techniques inherently carry dual-use risks. While watermarking enables provenance tracking to combat misuse, it could potentially be exploited for punitive enforcement based on probabilistic signals. Adversarial actors may further weaponize watermarking to falsely attribute content to specific organizations or models. We additionally acknowledge residual concerns regarding false positives, which could lead to wrongful accusations. We recommend deployment with conservative thresholds, transparent documentation of error rates, and appeals processes for affected users, calibrated to the specific deployment context.
@@
Acknowledgments
This work was supported in part by the National Science Foundation under Grant No. OAC-2112606. ChatGPT and Claude were used in the preparation of this work for editing and grammar checking. No passages were copied without full author review, and all substantive ideas, analysis, and conclusions are the product and responsibility of the authors. The AI tools were also utilized for code development and early manuscript reviews.
@@ Acknowledgments
      TBW Robustness ROC Curves ........................................................................................................................................................................E.7
@@
      False Positive Rate (FPR) on Human-Written Text ........................................................................................................................................................................E.8
@@
      Detection Attacks ........................................................................................................................................................................E.9
    • E.10
      Maximum $z$-Score Detection: Detailed Analysis ........................................................................................................................................................................E.10
@@ Appendix B Extended Threat Model
Excluded and Impractical Attacks. We further assume the adversary cannot mount impractical strategies that rely on excessive querying or access to paired datasets (watermarked vs. non-watermarked versions of the same prompt). Statistical comparison or supervised classification attacks are excluded from our primary model as they require assumptions that are rarely feasible for typical end-users. We provide an empirical evaluation of these “detection-based” attacks in E.9 for completeness, though they fall outside our primary security focus.
Partition Recovery Attacks. Prior work has demonstrated partition recovery attacks against KGW-style watermarks, where adversaries approximate green/red list assignments through repeated positional querying Jovanović et al. (2024). Because TBW uses a single fixed green list per generation rather than per-token hashing, this specific attack vector does not directly transfer. However, TBW’s semantic partitioning introduces a different concern: an adversary who knows the algorithm, topic list, and embedding approach might attempt to reconstruct the topic-aligned green lists through structural inference.
To evaluate this, we consider a strong attacker who knows TBW’s partitioning algorithm, the predefined topic set, and the general embedding approach, but not the private parameters: the specific embedding model weights, similarity threshold $\tau$, or the random seed used for round-robin distribution of residual tokens. We measure partition recovery across 12 attacker configurations, varying the embedding model (3 models, including an oracle configuration using the identical architecture as the watermark deployer) and similarity threshold ($\tau\in\{0.3,0.5,0.7,0.9\}$). For each configuration, we compute Jaccard similarity, precision, recall, and F1 between the attacker’s reconstructed partition and the true partition, averaged across all four topic lists. Tables 5 and 6 report results for OPT-6.7B and Gemma-7B, respectively.
Model $\boldsymbol{\tau}$ Jaccard Precision Recall F1
all-MiniLM 0.3 0.1433 0.2512 0.2514 0.2505
all-MiniLM 0.5 0.1448 0.2529 0.2529 0.2529
all-MiniLM 0.7 0.1419 0.2486 0.2486 0.2486
all-MiniLM 0.9 0.1415 0.2478 0.2478 0.2478
all-mpnet 0.3 0.1423 0.2491 0.2492 0.2491
all-mpnet 0.5 0.1414 0.2478 0.2478 0.2478
all-mpnet 0.7 0.1413 0.2476 0.2476 0.2476
all-mpnet 0.9 0.1415 0.2478 0.2478 0.2478
bge-small 0.3 0.1299 0.2546 0.2532 0.2256
bge-small 0.5 0.1301 0.2542 0.2532 0.2259
bge-small 0.7 0.1423 0.2491 0.2491 0.2491
bge-small 0.9 0.1438 0.2515 0.2515 0.2515
Table 5: Partition recovery attack results on OPT-6.7B across 12 attacker configurations. The all-MiniLM rows represent an oracle attacker using the identical embedding model as the watermark deployer.
Model $\boldsymbol{\tau}$ Jaccard Precision Recall F1
all-MiniLM 0.3 0.1427 0.2498 0.2498 0.2497
all-MiniLM 0.5 0.1422 0.2490 0.2490 0.2490
all-MiniLM 0.7 0.1425 0.2494 0.2494 0.2494
all-MiniLM 0.9 0.1434 0.2509 0.2509 0.2509
all-mpnet 0.3 0.1181 0.2489 0.2491 0.2064
all-mpnet 0.5 0.1428 0.2499 0.2499 0.2499
all-mpnet 0.7 0.1445 0.2525 0.2525 0.2525
all-mpnet 0.9 0.1434 0.2509 0.2509 0.2509
bge-small 0.3 0.0940 0.2502 0.2498 0.1617
bge-small 0.5 0.0940 0.2497 0.2497 0.1618
bge-small 0.7 0.1432 0.2506 0.2506 0.2506
bge-small 0.9 0.1431 0.2503 0.2503 0.2503
Table 6: Partition recovery attack results on Gemma-7B across 12 attacker configurations. Results are consistent with OPT-6.7B, confirming chance-level recovery across both models.
Even the oracle attacker using the identical embedding model (all-MiniLM) and matching $\tau$ achieves only chance-level recovery across both models. With $K=4$ topics, random assignment would yield Jaccard =0.25=0.25 and F1 =0.25=0.25 in expectation; the observed values fall at or below this baseline across all configurations. Some non-oracle configurations (e.g., bge-small at $\tau=0.3$) perform below chance, reflecting topic-specific misalignment that further underscores the difficulty of partition inference. This result stems from TBW’s dual assignment mechanism: tokens exceeding the similarity threshold $\tau$ are assigned semantically, while residual tokens are distributed via a randomly shuffled round-robin process (§4). Since the random seed governing the round-robin assignment is private, the full partition is unrecoverable regardless of the attacker’s knowledge of the embedding model or threshold. Semantic co-grouping alone is insufficient without the private round-robin seed. As complementary evidence, our word cloud analysis (Figure 9) confirms that watermarked outputs show no systematic elevation of topic-specific tokens, preventing observational inference of the active partition from generated text.